OpenStudy (anonymous):

Choose the Standard Form of the equation y =-1/4 x - 6. 1/4x + y = -6 4x + y = -6 x + 4y = -24 -4x + y = -24 Extremely confused:/

OpenStudy (diyadiya):

\[y=\frac{-1}{4}x-6\] \[\frac{1}{4}x+y=-6\] Multiply each term by 4 \[x + 4y = -24\]

OpenStudy (anonymous):

thank you !

OpenStudy (diyadiya):

:)